OpenStudy (anonymous):

Robby invested $18,400 at 7% interest. he earned at the end of x year what is the value of x ?

OpenStudy (anonymous):

ok i need help

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

what do i do

OpenStudy (anonymous):

help me plz

OpenStudy (ilovewolf):

its you answer love xd

OpenStudy (anonymous):

18400 * .07 = 1,288 9016 / 1288 = 7 1288 per year

OpenStudy (anonymous):

it is d

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

can you help me on the next ? plz

OpenStudy (anonymous):

I can try :) tag me

OpenStudy (anonymous):

Jacob invested $300 in saving account and earned $81 in interest at the end of 9 years. what was the interest rate

OpenStudy (anonymous):

so what is it

OpenStudy (ilovewolf):

http://www.weegy.com/?ConversationId=B80FEE22

OpenStudy (anonymous):

not you him

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

@ilovewolf um please stop putting in websites crvp cuz it never helps!!!

OpenStudy (anonymous):

yay it dot

OpenStudy (anonymous):

help me plz

OpenStudy (anonymous):

Formula for simple interest is: I = P*r*t where P is the Principal ($300) r is the interest rate in decimal (?) t is the number of years invested (9) and I is the interest earned ($81). Substitute the values into the formula and solve for r.

OpenStudy (anonymous):

\[$300\times$81\times9\]

OpenStudy (anonymous):

do that

OpenStudy (anonymous):

yes or no

OpenStudy (anonymous):

The "r" you solve will be in decimal. Multiply it by 100 to express it as a percentage.

OpenStudy (anonymous):

\[$300\times100\]

OpenStudy (anonymous):

that

OpenStudy (anonymous):

what are ur answer choices

OpenStudy (anonymous):

\[$300\times$100\times$81\]

OpenStudy (anonymous):

a is 3% and b is 3.5% and c is 4% and d is 4.5%

OpenStudy (anonymous):

can u take a good guess for that answer

OpenStudy (anonymous):

no can you give me the a plz

OpenStudy (anonymous):

it is a test 5 Q in the test

OpenStudy (anonymous):

if u take one good guess ill tell u the answer :P

OpenStudy (anonymous):

on the laset q5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

so what is the A

OpenStudy (anonymous):

so what do u think YOUR answer is

OpenStudy (anonymous):

can you give it to me

OpenStudy (anonymous):

im not gonna give u the answer away :P the person with the question need to take guess first then ill give u the answer

OpenStudy (anonymous):

i did that it is woung

OpenStudy (ilovewolf):

-_-

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

its 3% o-o